OpenStudy (anonymous):

The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle.

OpenStudy (anonymous):

"The sum of the measures of two complementary angles exceeds the difference of their measures by 72", so x+y = x-y + 72 y = -y + 72 y + y = 72 2y = 72 y = 72/2 = 36